James is building a ramp for his elderly grandmother. If he builds it 3 feet above the ground and if it has a 4 foot base, how long will the ramp be?
A. 3 feet
B. 5 feet
C. 8 feet
D. 10 feet

Answer by nerdybill(7384)

James is building a ramp for his elderly grandmother. If he builds it 3 feet above the ground and if it has a 4 foot base, how long will the ramp be?
A. 3 feet
B. 5 feet
C. 8 feet
D. 10 feet
.
It's a 3-4-5 triangle -- so it would be B.
Or, you can apply Pythagorean theorem:
Let r = ramp
then
r^2 = 3^2 + 4^2
r^2 = 9 + 16
r^2 = 25
r = 5 feet